DAY 02/01/00 shift summary


Subject: DAY 02/01/00 shift summary
From: Vivian O'Dell (odell@fnal.gov)
Date: Wed Jan 12 2000 - 16:27:54 CST


On Shift: SashaL

No beam. Running cosmics all the time without fruits (B09 is off). Etot
is off. Open access. JasonL is working on fruits. DC1Y (U and D) Raw Time
distribution on cosmics looks very strange. DC experts, please look at it.

Sydney was here trying to understand recent accident when data tape was
overwritten. He suggested the following test which was done after he left:

1. Fresh tape was inserted in ktev3/tps2d1
2. Cosmic run 15352 is written on this tape with OVERWRITE option.
3. run 15352 is stopped and 15353 started with APPEND option.
4. run 15353 is interrupted with drt_fresh_start while triggers are
   running.
5. run 15354 is initialized with APPEND option. We expected that
   initialization will hang at this point. However, run was initialized
   and triggers start running without any errors.
6. run 15354 is stopped. Tape is still in ktev3/tps2d1 and it is labelled
   as "Test 02/01/00".

The question is "What is on this tape?"

SashaL



This archive was generated by hypermail 2b27 : Wed Jan 12 2000 - 16:27:57 CST